# Frustrated Random Walks: A Fast Method to Compute Node Distances on Hypergraphs

### v2

@@ Abstract
A hypergraph is a generalization of a graph that arises naturally when attribute-sharing among entities is considered. Compared to graphs, hypergraphs have the distinct advantage that they contain explicit communities and are more convenient to manipulate. An open problem in hypergraph research is how to accurately and efficiently calculate node distances on hypergraphs. Estimating node distances enables us to find a node’s nearest neighbors, which has important applications in such areas as recommender system, targeted ads, etc. In this paper, we propose using expected hitting times of random walks to compute hypergraph node distances. We note that simple random walks (SRW) cannot accurately compute node distances on highly complex real-world hypergraphs, which motivates us to introduce frustrated random walks (FRW) for this task. We further benchmark our method against DeepWalk, and show that while the latter can achieve comparable results, FRW has a distinct computational advantage in cases where the number of targets is fairly small. For such cases, we show that FRW runs in significantly shorter time than DeepWalk. Finally, we analyze the time complexity of our method, and show that for large and sparse hypergraphs, the complexity is approximately linear, rendering it superior to the DeepWalk alternative.
@@ I Introduction
Hypergraphs arise naturally when there are explicit communities of nodes. In social media, for instance, users are represented as nodes, and users belonging to the same group are collected into the same hyperedge. With the hypergraph representation of users, we want to quantify the similarity between any pair of users, and use this information to suggest friends or recommend products. Intuitively, the similarity between users should be negatively correlated with their distance, thus the computation of hypergraph node distances is a prerequisite for the implementation of recommender systems and targeted adsSarwar et al. (2000); Fouss et al. (2007); Wu et al. (2022).
Although we can easily convert a hypergraph into a graph by expanding its hyperedges, we claim that hypergraphs contain more information than what can be encapsulated into its expanded graphs due to the explicit presence of communities in the former. Furthermore, the expansion of hyperedges can be computationally expensive and memory-intensive. It is thus advantageous to directly work with hypergraphs whenever possible rather than with their expanded graphs.
@@
Given the important applications of computing hypergraph node distances, we have developed a method based on random walks for this task. We will use the expected hitting times of random walks as node distances on hypergraphs. We will show that the simple random walk method is insufficient for the description of complex human-human interactions, and propose our frustrated random walk method. In frustrated random wallks, we augment each transition with an acceptance probability, and thus obtain a different transition matrix from the one in simple random walks. Although the simple and frustrated random walks have different transition matrices, we have developed a unified theoretical framework to efficiently compute their expected hitting times. In summary, the major contributions of our work are as follows:
  1. 1.
    We propose frustrated random walks on hypergraphs, a random walk scenario that outperforms simple random walks for computing node distances in heavily-weighted and scale-free hypergraphs.
  2. 2.
    We design a unified theoretical framework for computing the expected hitting times of random walks, whether it be simple or frustrated, and use the expected hitting times as hypergraph node distances.
  3. 3.
    We test our methods on real-world hypergraph datasets and demonstrate that frustrated random walk method significantly outperforms existing methods in accuracy and/or running speed.
@@ II Background
The importance of the computation of graph node distances has already been widely acknowledgedSommer et al. (2016), with significant applications in social networks (friend suggestion, link prediction, targeted ads), e-commerce (product recommendation), etc. Intuitively, the shorter the distance between two nodes, the more similar they are to each other. Therefore, the primary purpose of computing node distances is to determine node similarities, thus enabling the identification of the nearest neighbors of a target node within a graph. An obvious candidate for computing graph node distances is Dijkstra’s algorithm, which can quickly provide the shortest or geodesic distance from a source node to a target node. Although this algorithm is quick, it focuses solely on the shortest path and ignores all possible alternative routes. However, in a graph representation of social networks, the similarity between two nodes is expected to increase with the number of available alternative routes connecting them. This expectation has led to the development of algorithms that consider the entire graph structure when computing node distances. A plethora of such algorithms, which can be crudely classified into two categories, already exist in the literature. Below, we provide a brief review of these algorithms, focusing only on those most relevant to our work, rather than attempting an exhaustive review of all existing methods.
@@
The first category of algorithms relies on the study of random walks on graphs, and uses expected hitting time (also known as first-passage time) and commute time to evaluate node distancesKlein and Randić (1993); White and Smyth (2003); Fouss et al. (2007); Li and Le (2020). In this framework, for a pair of graph nodes denoted as $v_{i}$ and $v_{j}$, the hitting time $N_{j}^{(i)}$ of a random walk process is defined as the number of steps a random walker starting from node $v_{i}$ needs to traverse before it hits node $v_{j}$ for the first time. $N_{j}^{(i)}$ is a random variable whose expectation value $\mathbb{E}N_{j}^{(i)}$ can be used to measure the distance from $v_{i}$ to $v_{j}$. Here, the distance should be understood as the dissimilarity between nodes $v_{i}$ and $v_{j}$ since it does not satisfy the symmetricity and triangle-inequality conditions that are required of a mathematical distance function. From the expected hitting times, we can construct the expected commute time, defined as $n(i,j)=\mathbb{E}N_{j}^{(i)}+\mathbb{E}N_{i}^{(j)}$, which satisfies all the conditions of a distance function in mathematics, and is therefore a genuine distance in the strict mathematical sense. Given a graph structure, the closed form expressions for $\mathbb{E}N_{j}^{(i)}$ and $n(i,j)$ can be readily obtained from the pseudo-inverse of the graph Laplacian, as is shown in Ref. Fouss et al. (2007). It is however shown in Ref. Li and Le (2020) that for complex and scale-free graphs whose node degree distribution follows a power law, the expected hitting times of simple random walks cannot accurately evaluate node distances, which necessitates the introduction of frustrated random walks to graphs. Here, we will generalize these methods to hypergraphs, and show that frustrated random walks have distinct advantage in measuring node distances of heavily weighted and scale-free hypergraphs.
@@
The second category of algorithms relies on graph node embedding, also called graph representation learning, which means mapping graph nodes to low-dimensional dense vectors and using the cosine distances between these vectors as graph node distancesHamilton et al. (2017a). This includes such algorithms as DeepWalkPerozzi et al. (2014), node2vecGrover and Leskovec (2016), structural deep network embedding (SDNE)Wang et al. (2016), graphSAGEHamilton et al. (2017b), graph attention networksVeličković et al. (2017), etc. These algorithms aim to map graph nodes to a continuous vector space so that nodes that are neighbors in the original graph are mapped to vectors with small cosine distances. There is no universal definition of neighborhood in these algorithms. For example, in DeepWalk, node2vec and graphSAGE, nodes that frequently co-appear in sampled random paths are considered neighbors, whereas in SDNE, neighboring nodes are those that are of first-order or second-order proximity with respect to each other in the original graph. Despite these differences, all algorithms adhere to the guiding principle that neighboring nodes should be mapped to neighboring vectors during the node embedding process. After node embedding, we can use cosine distances between mapped node vectors as graph node distances. Generally, the cosine distances from the node embedding methods is a good metric for evaluating graph node distances, however complex the graphs may be. Among these algorithms, the DeepWalk method excels at measuring node distances of scale-free graphs. A second advantage of DeepWalk in comparison with the other node embedding algorithms is the ease with which it can be generalized to hypergraphs, which prompts us to use it as our benchmark method in our work.
@@
Although we can easily generalize DeepWalk to hypergraphs, for the other algorithms, such a generalization is highly non-trivial. In this paper, we generalize frustrated random walks to hypergraphs, and show that its performance is on par with the that of DeepWalk even for complex and scale-free hypergraphs. Compared to DeepWalk, frustrated random walks (FRW) has the following four advantages. First, in applications where the goal is to find the nearest neighbors of a few nodes in a large hypergraph, FRW is a preferable and faster option than DeepWalk. Second, FRW conveniently gives a closed-form and interpretable solution for node distances, a solution that is impossible to obtain using traditional deep-learning-based methods. Third, FRW does not require any parameter tuning, whereas DeepWalk requires heavy investment in this task. Fourth, and unlike that in DeepWalk or any other node embedding algorithm, the node distances of FRW are asymmetric, rendering it more suitable to describe real-world relationships which are generally non-equivalent and non-reflective.
@@ IV.1 Hypergraphs with ground-truth node labels
Figure 2: Log-log plot of node and hyperedge degree distribution for arXiv data set. Although the node degree distribution does not strictly follows the power law, the power law is pretty obvious for hyperedge degree distribution.
@@
The second data set we experiment with is the trivago-clicks data setBenson (2021). This data set has been introduced by the authors of Ref. Chodrow et al. (2021) as a collection of behavior data from users browsing trivago.com while trying to book a hotel. In this hypergraph data set, nodes represent accommodations (mostly hotels) browsed by users, and hyperedges represent a user’s browsing history in the same browsing session before checkout (i.e., before an order is placed). Each node (accommodation) is associated to a country, which can be used as the node’s label. The data set contains $172,738$ nodes and $233,202$ hyperedges. The number of distinct node labels (countries or regions) is $160$. The node degree follows a power law, yet the node degrees are generally small, ranging from $1$ to $339$, which is significantly smaller than the hypergraph node number. As a result, this hypergraph is still lightly weighted. A log-log plot of the node degree distribution is shown in Fig. 3.
Figure 3: Log-log plot of node degree distribution for trivago data set. This data set is a scale-free hypergraph.
@@
In this section, we have shown the effectiveness of FRW in measuring node distances on real-world hypergraphs. Although FRW is effective, its advantage compared to SRW is not yet fully visible. In this next section, we will demonstrate the full advantage of FRW using more heavily-weighted and scale-free hypergraphs.
@@ IV.2 Heavily-weighted and scale-free hypergraphs without node labels
As already stated in the previous sections, the node degree distribution of a scale-free hypergraph that is generated according to the principle of preferential attachment follows the power lawBarabási and Albert (1999). A hypergraph is called heavily weighted if its node degree range spans several orders of magnitude, and the maximum node degree is at least as large as the hypergraph node number. Since we have designed the frustrated random walks (FRW) with preferential attachment in mind, we expect this method to more accurately compute node distances on scale-free hypergraphs, as will be demonstrated in this section.
@@
We first apply our method to a hypergraph constructed from the seven Harry Potter novels written by J.K. Rowling. This novel was chosen for its intricate plots, complex character relationships and world-wide popularity. In our hypergraph, characters are represented as nodes, and each scene in the novel is represented as a hyperedge. Characters are grouped into the same hyperedge if they appear in the same scene. Node degree indicates the number of times a node (character) appears in the novel. This hypergraph contains $183$ nodes and $3557$ hyperedges. The maximum node degree is 4548, which is much larger than node number. Thus, this hypergraph is heavily-weighted. In this hypergraph, the node number is too small for the power law to be obviously visible. One way to see power law in this data set is to expand the hyperedges into edges, and plot the log-log distribution of the edge weights, as shown in Fig. 4.
Figure 4: Log-log plot of edge weight distribution in Harry Potter data set. This figure shows the data set is a heavily-weighted, scale-free hypergraph.
From our knowledge of the novels, we know that Ron Weasley and Hermione Granger are Harry Potter’s best friends. Thus, a good node-distance-computing algorithm should find these two characters as Harry Potter’s top-2 nearest neighbors, meaning their distances to Harry Potter should be the smallest among all possible nodes. This is the golden criteria we employ in this section to check the validity of our algorithms. We use SRW, FRW, and DeepWalk to rank the nodes by their distances from near to far with respect to Harry Potter, and list our results in Table 4. From the table, we can see that both FRW and DeepWalk list Ron Weasley and Hermione Granger as the nearest neighbors to Harry Potter, yet SRW lists Hokey as the nearest neighbor, which is in sharp contrast with our intuition. This corroborates our assertion that FRW is more suitable than SRW for describing real-world human relationships.
@@
We then apply our results to a hypergraph created from Dream of the Red Chamber, a novel written by Cao Xueqin. This novel is also famous for its complicated plots and tangled character relationships. Still, we represent characters using nodes and characters appearing in the same scene are grouped into the same hyperedge. This hypergraph contains $324$ nodes and $1,524$ hyperedges. The node degree distribution follows a power law, as shown in Fig 5. The node degrees range from $1$ to $954$ which is much larger than the node number, meaning this hypergraph is also heavily weighted. We will show that for such a hypergraph, the FRW method is most suitable for calculating node distances.
Figure 5: Log-log plot of node degree distribution in Dream of the Red Chamber. This figure shows the data set is a heavily-weighted, scale-free hypergraph.
We will again use DeepWalk as our benchmark, and use SRW and FRW to obtain the nearest neighbors for each target node. The Spearman correlation coefficients between SRW/FRW results and DeepWalk results quantify the accuracy of ranking results from SRW and FRW. In Table 7, we show the Spearman correlation coefficients for nodes of various degrees. From this table, we can reach a similar conclusion as that with Harry Potter data set. That is, for central nodes with high degrees, FRW and DeepWalk can yield results that meet our intuition, whereas SRW deviates significantly from either FRW or DeepWalk. Only for peripheral nodes of low degrees can the performance of SRW be on par with that of the other two methods.
@@
V Time complexity of the expected hitting time method
$(I-B)\boldsymbol{x}=\mathbbm{1}$ (17)
For large and sparse coefficient matrix $B$, we can only solve this equation using conjugate gradient method or its variants. Conjugate gradient method involves a series of iteration cycles, and the most time consuming part of each cycle is the evaluation of $Bv$, where $B$ is the probability transition matrix and $v$ is a dense vector. Thus, the time complexity of our method as described in Algorithm 1 is $N_{iter}\times O(Bv)$, where $N_{iter}$ is the iteration number in conjugate gradient method and $O(Bv)$ is the time complexity of evaluating $Bv$. We will determine $O(Bv)$ first.
VI Comparison of running speeds
In the previous sections, we have shown that to find a node’s nearest neighbors using SRW/FRW methods, we only need to construct the transition matrix from the hypergraph structure and the target node, and solve a system of linear equations using conjugate gradient method, the time complexity of which is approximately linear. However, if we want to accomplish this using DeepWalk, we need to map all the nodes to vectors, calculate the cosine distances between the target node vector and all the other node vectors, and finally rank all the other nodes with respect to the target according to cosine distances. The node mapping could be pretty time consuming, thus rendering the DeepWalk inferior in speed compared to SRW/FRW methods.

### v3

@@ Abstract
A hypergraph is a generalization of a graph that arises naturally when attribute-sharing among entities is considered. Compared to graphs, hypergraphs have the distinct advantage that they contain explicit communities and are more convenient to manipulate. An open problem in hypergraph research is how to accurately and efficiently calculate node distances on hypergraphs. Estimating node distances enables us to find a node’s nearest neighbors, which has important applications in such areas as recommender system, targeted advertising, etc. In this paper, we propose using expected hitting times of random walks to compute hypergraph node distances. We note that simple random walks (SRW) cannot accurately compute node distances on highly complex real-world hypergraphs, which motivates us to introduce frustrated random walks (FRW) for this task. We further benchmark our method against DeepWalk, and show that while the latter can achieve comparable results, FRW has a distinct computational advantage in cases where the number of targets is fairly small. For such cases, we show that FRW runs in significantly shorter time than DeepWalk. Finally, we analyze the time complexity of our method, and show that for large and sparse hypergraphs, the complexity is approximately linear, rendering it superior to the DeepWalk alternative.
@@ I Introduction
Hypergraphs arise naturally when there are explicit communities of nodes. In social media, for instance, users are represented as nodes, and users belonging to the same group are collected into the same hyperedge. With the hypergraph representation of users, we want to quantify the similarity between any pair of users, and use this information to suggest friends or recommend products. Intuitively, the similarity between users should be negatively correlated with their distance, thus the computation of hypergraph node distances is a prerequisite for the implementation of recommender systems and targeted advertisingSarwar et al. (2000); Fouss et al. (2007); Wu et al. (2022).
Although we can easily convert a hypergraph into a graph by expanding its hyperedges, we claim that hypergraphs contain more information than what can be encapsulated into its expanded graphs due to the explicit presence of communities in the former. Furthermore, the expansion of hyperedges can be computationally expensive and memory-intensive, thus making it advantageous to directly work with hypergraphs whenever possible rather than with their expanded graphs.
@@
Given the important applications of computing hypergraph node distances, we have developed a method based on random walks for this task. We will use the expected hitting times of random walks as node distances on hypergraphs. We will show that the simple random walk method is insufficient for the description of complex human-human interactions, and propose our frustrated random walk method. In frustrated random wallks, we augment each transition with an acceptance probability, and thus obtain a different transition matrix from the one in simple random walks. Although the simple and frustrated random walks have different transition matrices, we have developed a unified theoretical framework to efficiently compute their expected hitting times.
In summary, we made three major contributions in our work. First, we propose frustrated random walks on hypergraphs, a random walk scenario that outperforms simple random walks in computing node distances for heavily-weighted and scale-free hypergraphs. Second, we design a unified theoretical framework for computing the expected hitting times of random walks, whether it be simple or frustrated, and use the expected hitting times as hypergraph node distances. Third, we test our methods on real-world hypergraph datasets and demonstrate that frustrated random walk method significantly outperforms existing methods in accuracy and/or running speed.
@@ II Background
The importance of the computation of graph node distances has already been widely acknowledgedSommer et al. (2016), with significant applications in social networks (friend suggestion, link prediction, targeted advertising), e-commerce (product recommendation), etc. Intuitively, the shorter the distance between two nodes, the more similar they are to each other. Therefore, the primary purpose of computing node distances is to determine node similarities, thus enabling the identification of the nearest neighbors of a target node within a graph. An obvious candidate for computing graph node distances is Dijkstra’s algorithm, which can quickly provide the shortest or geodesic distance from a source node to a target node. Although this algorithm is quick, it focuses solely on the shortest path and ignores all possible alternative routes. However, in a graph representation of social networks, the similarity between two nodes is expected to increase with the number of available alternative routes connecting them. This expectation has led to the development of algorithms that consider the entire graph structure when computing node distances. A plethora of such algorithms already exist in the literature, which can be crudely classified into two categories. Below, we provide a brief review of these algorithms, focusing only on those most relevant to our work, rather than attempting an exhaustive review of all existing methods.
@@
The first category of algorithms relies on the study of random walks on graphs, and uses expected hitting time (also known as first-passage time) and commute time to evaluate node distancesKlein and Randić (1993); White and Smyth (2003); Fouss et al. (2007); Li and Le (2020). In this framework, for a pair of graph nodes denoted as $v_{i}$ and $v_{j}$, the hitting time $N_{j}^{(i)}$ of a random walk process is defined as the number of steps a random walker starting from node $v_{i}$ needs to traverse before it hits node $v_{j}$ for the first time. $N_{j}^{(i)}$ is a random variable whose expectation value $\mathbb{E}N_{j}^{(i)}$ can be used to measure the distance from $v_{i}$ to $v_{j}$. Here, the distance should be understood as the dissimilarity between nodes $v_{i}$ and $v_{j}$ since it does not satisfy the symmetricity and triangle-inequality conditions that are required of a mathematical distance function. From the expected hitting times, we can construct the expected commute time, defined as $n(i,j)=\mathbb{E}N_{j}^{(i)}+\mathbb{E}N_{i}^{(j)}$, which satisfies all the conditions of a distance function in mathematics, and is therefore a genuine distance in the strict mathematical sense. Given a graph structure, the closed form expressions for $\mathbb{E}N_{j}^{(i)}$ and $n(i,j)$ can be readily obtained from the pseudo-inverse of the graph Laplacian, as is shown in Ref. Fouss et al. (2007). It is however shown in Ref. Li and Le (2020) that for heavily-weighted and scale-free graphs whose node degree distribution follows a power law, the expected hitting times of simple random walks cannot accurately evaluate node distances, which necessitates the introduction of frustrated random walks to graphs. Here, we will generalize these methods to hypergraphs, and show that frustrated random walks have distinct advantages in measuring node distances of such highly complicated hypergraphs.
@@
The second category of algorithms relies on graph node embedding, also called graph representation learning, which means mapping graph nodes to low-dimensional dense vectors and using the cosine distances between these vectors as graph node distancesHamilton et al. (2017a). This includes such algorithms as DeepWalkPerozzi et al. (2014), node2vecGrover and Leskovec (2016), structural deep network embedding (SDNE)Wang et al. (2016), graphSAGEHamilton et al. (2017b), graph attention networksVeličković et al. (2017), etc. These algorithms aim to map graph nodes to a continuous vector space so that nodes that are neighbors in the original graph are mapped to vectors with small cosine distances. There is no universal definition of neighborhood in these algorithms. For example, in DeepWalk, node2vec and graphSAGE, nodes that frequently co-appear within a pre-specified window in sampled random paths are considered neighbors, whereas in SDNE, neighboring nodes are those that are of first-order or second-order proximity with respect to each other in the original graph. Despite these differences, all algorithms adhere to the guiding principle that neighboring nodes should be mapped to neighboring vectors during the node embedding process. After node embedding, we can use cosine distances between mapped node vectors as graph node distances. Generally, the cosine distances from the node embedding methods is a good metric for evaluating graph node distances, however complex the graphs may be. Among these algorithms, the DeepWalk method excels at measuring node distances of scale-free graphs. A second advantage of DeepWalk in comparison with the other node embedding algorithms is the ease with which it can be generalized to hypergraphs, which prompts us to use it as our benchmark method in our work.
@@
Although we can easily generalize DeepWalk to hypergraphs, for the other algorithms, such a generalization is highly non-trivial. In this paper, we generalize frustrated random walks to hypergraphs, and show that its performance is on par with the that of DeepWalk even for highly complicated hypergraphs. Compared to DeepWalk, frustrated random walks (FRW) has the following four advantages. First, in applications where the goal is to find the nearest neighbors of a few nodes in a large hypergraph, FRW is a preferable and faster option than DeepWalk. Second, FRW conveniently gives a closed-form and interpretable solution for node distances, a solution that is impossible to obtain using traditional deep-learning-based methods. Third, FRW does not require any parameter tuning, whereas DeepWalk requires heavy investment in this task. Fourth, and unlike that in DeepWalk or any other node embedding algorithm, the node distances of FRW are asymmetric, rendering it more suitable to describe real-world relationships which are generally non-equivalent and non-reflective.
@@ IV.1 Hypergraphs with ground-truth node labels
Figure 2: $\log_{10}-\log_{10}$ plot of node and hyperedge degree distribution for arXiv data set. Although the node degree distribution does not strictly follows the power law, the power law is pretty obvious for hyperedge degree distribution.
@@
The second data set we experiment with is the trivago-clicks data setBenson (2021). This data set has been introduced by the authors of Ref. Chodrow et al. (2021) as a collection of behavior data from users browsing trivago.com while trying to book a hotel. In this hypergraph data set, nodes represent accommodations (mostly hotels) browsed by users, and hyperedges represent a user’s browsing history in the same browsing session before checkout (i.e., before an order is placed). Each node (accommodation) is associated to a country, which can be used as the node’s label. The data set contains $172,738$ nodes and $233,202$ hyperedges. The number of distinct node labels (countries or regions) is $160$. The node degree follows a power law, yet the node degrees are generally small, ranging from $1$ to $339$, which is significantly smaller than the hypergraph node number. As a result, this hypergraph is still lightly weighted. A $\log_{10}-\log_{10}$ plot of the node degree distribution is shown in Fig. 3.
Figure 3: $\log_{10}-\log_{10}$ plot of node degree distribution for trivago data set. This data set is a scale-free hypergraph.
@@
In this section, we have shown the effectiveness of FRW in measuring node distances on real-world hypergraphs. Although FRW is effective, its advantage compared to SRW is not yet fully visible. In the next section, we will demonstrate the full advantage of FRW using more heavily-weighted and scale-free hypergraphs.
@@ IV.2 Heavily-weighted and scale-free hypergraphs without node labels
As already stated in the previous sections, the node degree distribution of a scale-free hypergraph that is generated according to the principle of preferential attachment follows the power lawBarabási and Albert (1999). A hypergraph is called heavily weighted if its node degree range spans several orders of magnitude, and the maximum node degree is at least as large as the hypergraph node number. Since we have designed the frustrated random walks (FRW) with preferential attachment effect in mind, we expect this method to more accurately compute node distances on scale-free hypergraphs, as will be demonstrated in this section.
@@
We first apply our method to a hypergraph constructed from the Harry Potter novels written by J.K. Rowling. This novel was chosen for its intricate plots, complex character relationships and world-wide popularity. In our hypergraph, characters are represented as nodes, and each scene in the novel is represented as a hyperedge. Characters are grouped into the same hyperedge if they appear in the same scene. Node degree indicates the number of times a node (character) appears in the novel. This hypergraph contains $183$ nodes and $3557$ hyperedges. The maximum node degree is 4548, which is much larger than node number. Thus, this hypergraph is heavily-weighted. In this hypergraph, the node number is too small for the power law to be obviously visible. One way to see power law in this data set is to expand the hyperedges into edges, and plot the $\log_{10}-\log_{10}$ distribution of the edge weights, as shown in Fig. 4.
Figure 4: $\log_{10}-\log_{10}$ plot of edge weight distribution in Harry Potter data set. This figure shows the data set is a heavily-weighted, scale-free hypergraph.
From our knowledge of the novels, we know that Ron Weasley and Hermione Granger are Harry Potter’s best friends. Thus, a good node-distance-computing algorithm should find these two characters as Harry Potter’s top two nearest neighbors, meaning their distances to Harry Potter should be the smallest among all nodes. This is the golden criteria we employ in this section to check the validity of our algorithms. We use SRW, FRW, and DeepWalk to rank the nodes by their distances from near to far with respect to Harry Potter, and list our results in Table 4. From the table, we can see that both FRW and DeepWalk list Ron Weasley and Hermione Granger as the nearest neighbors to Harry Potter, yet SRW lists Hokey as the nearest neighbor, which is in sharp contrast with our intuition. This corroborates our assertion that FRW is more suitable than SRW for describing real-world human relationships.
@@
We then apply our methods to a hypergraph created from Dream of the Red Chamber, a novel written by Cao Xueqin. This novel is also famous for its complicated plots and tangled character relationships. Still, we represent characters using nodes and characters appearing in the same scene are grouped into the same hyperedge. This hypergraph contains $324$ nodes and $1,524$ hyperedges. The node degree distribution follows a power law, as shown in Fig 5. The node degrees range from $1$ to $954$ which is much larger than the node number, meaning this hypergraph is also heavily weighted. We will show that for such a hypergraph, the FRW method is most suitable for calculating node distances.
Figure 5: $\log_{10}-\log_{10}$ plot of node degree distribution in Dream of the Red Chamber. This figure shows the data set is a heavily-weighted, scale-free hypergraph.
To see this, we will again use DeepWalk as our benchmark, and use SRW and FRW to obtain the nearest neighbors for each target node. The Spearman correlation coefficients between SRW/FRW results and DeepWalk results quantify the accuracy of ranking results from SRW and FRW. In Table 7, we show the Spearman correlation coefficients for nodes of various degrees. From this table, we can reach a similar conclusion as that with Harry Potter data set. That is, for central nodes with high degrees, FRW and DeepWalk can yield results that meet our intuition, whereas SRW deviates significantly from either FRW or DeepWalk. Only for peripheral nodes of low degrees can the performance of SRW be on par with that of the other two methods.
@@
V Comparison of running speeds
In the previous sections, we have shown that to find a node’s nearest neighbors using SRW/FRW methods, we only need to construct the transition matrix from the hypergraph structure and the target node, and solve a system of linear equations using conjugate gradient method, the time complexity of which is approximately linear, as we will demonstrate in the next section. However, if we want to accomplish this using DeepWalk, we need to map all the nodes to vectors, calculate the cosine distances between the target node vector and all the other node vectors, and finally rank all the other nodes with respect to the target according to cosine distances. The node mapping could be pretty time consuming, thus rendering the DeepWalk inferior in speed compared to SRW/FRW methods.